OpenStudy (anonymous):

Inverse functions? The functions are: sin^-1(.65) tan^-1(-.09) I have to put them in radians

OpenStudy (anonymous):

Where is the function ?

OpenStudy (anonymous):

@Noura11 i just edited the question

OpenStudy (anonymous):

In this case you should use the calculator ! \[\sin^{-1}(0.65)=0.707\text{ rd}\]

OpenStudy (anonymous):

my options are these: http://puu.sh/3cR2l.png that's why i got confused

OpenStudy (anonymous):

and these are the options for \[\tan^{-1} (-.09)\] http://puu.sh/3cR9Q.png

OpenStudy (anonymous):

Yes ! In general case : to find \[\sin^{-1}(\alpha)=\beta +2k\pi\text{ or } (\pi-\beta)+2k\pi~~k\in{\mathbb Z}\] where \[\beta\] is the number that we will get by using the calculator

OpenStudy (anonymous):

okay, the answer i got from the calculator was like you said, .707 or .71. so it'd be .71 + 2kpi?

OpenStudy (anonymous):

The answer should be like this : \[\sin^{-1}(0.65)=.71+2k\pi \text{ or } (\pi-.71)+2k\pi\]

OpenStudy (anonymous):

but there is no option for 2kpi. would it be the same as 2pi n?

OpenStudy (anonymous):

Yes ! k or n ! There is no difference !

OpenStudy (anonymous):

okay, and the one for tan^-1(-.09) is option a?

OpenStudy (anonymous):

NO ! It is tha last one ! Because we have : \[\tan^{-1}(\alpha)=\beta +n\pi\] where \[\beta\] is the number that we will get by using the calculator !